[OH-] = 5.2 x 10 - 8
1
Expert's answer
2017-05-13T03:35:09-0400
A pH value is equal to -lg[H+] or 14 + lg[OH-]. Thus, pH = 14 - 8 + 0.72 = 6.72 (slightly acidic environment).